OpenStudy (anonymous):

What is the solution of the proportion 5/x+4=3/x-2?

OpenStudy (phi):

one thing you could do is flip both of them first.

OpenStudy (phi):

or just cross multiply

OpenStudy (anonymous):

what do you mean by fliping it

OpenStudy (phi):

if you have \[ \frac{5}{x+4} = \frac{3}{x-2} \] you could write it as \[ \frac{x+4}{5} = \frac{x-2}{3} \] maybe this way you can see how to solve it? I would either cross multiply or multiply both sides by 15 to get rid of the denominators.

OpenStudy (anonymous):

oh ok thx

OpenStudy (phi):

cross multiplying looks easy: \[ 5(x-2)= 3(x+4)\] distribute the 5 on the left side (multiply x and -2 by 5) distribute the 3 on the right side what do you get?

OpenStudy (anonymous):

yes i do but i tried that way n it did not work

OpenStudy (phi):

5(x−2)=3(x+4) what do you get after distributing the 5 and the 3?

OpenStudy (anonymous):

oh i earsed it but i got x = 7

OpenStudy (phi):

5(x−2)=3(x+4) what do you get after distributing the 5 and the 3?

OpenStudy (anonymous):

the flipping way i got x = 11